# Two-step estimation of latent trait models

### v2

@@ Abstract
We consider two-step estimation of latent variable models, in which just the measurement model is estimated in the first step and the measurement parameters are then fixed at their estimated values in the second step where the structural model is estimated. We show how this approach can be implemented for latent trait models (item response theory models) where the latent variables are continuous and their measurement indicators are categorical variables. The properties of two-step estimators are examined using simulation studies and applied examples. They perform well, and have attractive practical and conceptual properties compared to the alternative one-step and three-step approaches. These results are in line with previous findings for other families of latent variable models. This provides strong evidence that two-step estimation is a flexible and useful general method of estimation for different types of latent variable models.
Key words: Item response theory models; latent variable models; structural equation models; pseudo-maximum likelihood estimation
@@ 1 Introduction
Latent variable models have normally two main parts: a measurement model which describes how the latent variables that appear in the model are measured by observed indicators of them, and a structural model which describes the associations among the latent variables and any observed explanatory and response variables which are not treated as measurement indicators. For instance, in the illustrative example considered in Section 5 of this paper the structural model specifies how individuals’ extrinsic and intrinsic work value orientations are associated with characteristics of the individuals, and the measurement model specifies how these value orientations are measured by a set of survey questions. Here, as in many applications, the structural model is the focus of substantive interest, but the measurement model also needs to be included and estimated in order for the structural model to be estimable.
Estimation of these elements can be organised in different ways. In joint or one-step estimation, both parts of the model are estimated together. When this is done by maximizing the joint likelihood, one-step estimates are the maximum likelihood (ML) estimates of the model parameters. In contrast, “stepwise” approaches divide the estimation into separate steps. The most familiar of them is three-step estimation. In its first step, the measurement model is estimated from a specification which omits all or most of the structural model. In the second step, this estimated measurement model is used to assign predicted values of the latent variables to the units of analysis, such as factor scores for continuous latent variables. In the third step, these scores are used in place of the latent variables to estimate the structural model. There are “naive” and “adjusted” versions of three-step estimation, depending on whether it attempts to account for the measurement error that results from using the predicted values.
@@
Two-step estimation is a general idea, and one of the points that we want to make in this paper is that it can in principle be applied to any latent variable models. However, there may be differences in performance of the estimates, ease of implementation, or other points to consider when it is used for different broad classes of models. In the literature so far, two-step estimation has accordingly been described separately for different types of models. It was proposed for latent class analysis, that is for models where both the latent variables and their indicators are categorical, by Bandeen-Roche et al. (1997), Xue and Bandeen-Roche (2002) and Bakk and Kuha (2018), and has since been extended to further versions of them (e.g. multilevel latent class analysis in Di Mari et al. 2023). For structural equation models (SEMs), where the latent variables and indicators are both continuous, the idea was introduced already by Burt (1973, 1976), but detailed exploration of it is much more recent. In particular, Rosseel and Loh (2022) proposed a general two-step approach for SEMs, with the title of “structural-after-measurement” (SAM) estimation. They describe two (usually equivalent) forms of it: “local SAM” which summarises the first step in the form of the estimated means and covariance matrix of the variables in the structural model, and “global” SAM which is analogous to the implementation that we describe in this paper for latent trait models (the moment-based “local” approach is not really applicable to them). A method for SEMs is also proposed by Levy (2023), using Bayesian (MCMC) estimation for both steps. Another closely related paper is Skrondal and Kuha (2012), who use two-step estimation to correct for covariate measurement error in regression models. Two-step methods for these classes of models are now also being implemented in general-purpose software for latent variable modelling, in Latent Gold for latent class analysis (Vermunt and Magidson, 2021b) and in the R package lavaan for SAM (Rosseel, 2012; Rosseel and Loh, 2022).
@@
In this paper we propose and examine two-step estimation for another family of latent variable models, one where the latent variables are continuous but their indicators are categorical. These are commonly known as Item Response Theory (IRT) models. We refer to them also as latent trait models. They resemble SEMs in having continuous latent variables, and standard latent class models in having categorical indicators. Like SEMs, they are often used with complex structural models which include multiple latent variables, whereas latent class analysis more often includes only one latent class variable. Computationally, latent trait models are typically the most demanding of these families of models, at least for ML estimation, because their log likelihood cannot be expressed in a closed form.
@@
The latent trait model specifications that we consider are introduced in Section 2. The definition, implementation and properties of two-step estimation for them are described in Section 3. Simulation studies are reported in Section 4 and the real-data example in Section 5, and concluding discussion is given in Section 6. Extended tables of some of the results from the simulations and data analysis are given in a supplementary appendix. Computer code for the estimation in R and Mplus (both of which are needed) and replication code for the applied example are also provided as supplementary materials.
@@ 2 Models and variables
In two-step estimation, $\boldsymbol{\theta}_{1}$ are estimated first and then fixed at their estimated values in the second step where $\boldsymbol{\theta}_{2}$ are estimated. This is a very general approach which can in principle be used for any instance of this general model. For specificity, however, in this paper we focus on its use for a particular class of latent variable models. Here the items in $\mathbf{Y}_{i}$ are categorical variables, $\boldsymbol{\eta}_{i}$ are continuous and normally distributed, the structural model is recursive, and the measurement models exclude direct effects from $\mathbf{X}_{i}$ to $\mathbf{Y}_{i}$. This specification is described in this section. Some further variants and extensions of it are discussed in Section 3.5.
@@
where $\mathbf{B}_{0}$, $\mathbf{B}_{\eta}$ and $\mathbf{B}_{x}$ are matrices of parameters, and $\boldsymbol{\zeta}_{i}\sim N(\mathbf{0},\boldsymbol{\Psi})$ are normally distributed model residuals. This is a variant of a formulation that is commonly used to specify structural equation models in a compact and general form. For our purposes, however, it is more convenient to write the model in a different way, in terms of a chain of conditional distributions. This is possible because we consider only recursive models, obtained from (1) by imposing appropriate constraints on $\mathbf{B}_{\eta}$. Suppose that $\boldsymbol{\eta}_{i}=(\eta_{i1},\dots,\eta_{iK})^{\prime}$ includes $K\geq 1$ variables. Suppose further that the model specification groups them into $1\leq M\leq K$ blocks of variables $\boldsymbol{\eta}_{i(m)}$, so that we can also write $\boldsymbol{\eta}_{i}=(\boldsymbol{\eta}_{i(1)}^{\prime},\dots,\boldsymbol{\eta}_{i(M)}^{\prime})^{\prime}$ in this pre-specified order. The structural model can then be written as
@@
××$\hskip 10.00002pt\times\cdots\times p(\boldsymbol{\eta}_{i(M)}|\mathbf{X}_{i},\boldsymbol{\eta}_{i(1)},\dots,\boldsymbol{\eta}_{i(M-1)};\boldsymbol{\theta}_{2M})$ (2)
where the structural parameters are partitioned correspondingly as $\boldsymbol{\theta}_{2}=(\boldsymbol{\theta}_{21}^{\prime},\dots,\boldsymbol{\theta}_{2M}^{\prime})^{\prime}$. It follows from (1) that all of these conditional distributions are (multivariate or univariate) normal and specified by the linear models
@@
$p(\boldsymbol{\eta}_{i(m)}|\mathbf{X}_{i},\boldsymbol{\eta}_{i(1)},\dots,\boldsymbol{\eta}_{i(m-1)};\boldsymbol{\theta}_{2m})$ $\sim$ $N(\boldsymbol{\beta}_{m0}+\boldsymbol{\beta}_{mx}\mathbf{X}_{i}+\sum_{l=1}^{m-1}\boldsymbol{\beta}_{ml}\boldsymbol{\eta}_{i(l)},\,\boldsymbol{\Psi}_{m})$ (4)
for $m=2,\dots,M$, so that each of $\boldsymbol{\theta}_{21},\dots,\boldsymbol{\theta}_{2M}$ consists of the corresponding $\boldsymbol{\beta}$ and $\boldsymbol{\Psi}$ parameters. Constraints on elements of $\boldsymbol{\theta}_{2}$ can be included, most often zero constraints on the $\boldsymbol{\beta}$ parameters which omit some of the explanatory variables from some parts of the model and thus yield different non-saturated choices for the overall structural model (2). If observed covariates $\mathbf{X}_{i}$ are not present, they are omitted from (2)–(4). If any $\boldsymbol{\eta}_{i(m)}$ are actually observed (rather than latent) variables, it would also be straightfoward to replace the linear model with some other model for them. For example, if $\boldsymbol{\eta}_{i(M)}$ is a single binary observed variable, we could specify a binary logistic model for it given $\mathbf{X}_{i}$ and $\boldsymbol{\eta}_{i(1)},\dots,\boldsymbol{\eta}_{i(M-1)}$.
Consider then the measurement model. We assume throughout that $p(\mathbf{Y}_{i}|\boldsymbol{\eta}_{i},\mathbf{X}_{i};\boldsymbol{\theta}_{1})=p(\mathbf{Y}_{i}|\boldsymbol{\eta}_{i};\boldsymbol{\theta}_{1})$, so that there is no non-equivalence of measurement (differential item functioning) with respect to observed covariates $\mathbf{X}_{i}$. The overall model can then be written as
@@
[…]
@@
so that $\boldsymbol{\theta}_{1}=(\boldsymbol{\theta}_{11}^{\prime},\dots,\boldsymbol{\theta}_{1K}^{\prime})^{\prime}$, with the different $\boldsymbol{\theta}_{1k}$ taken to be distinct from each other. We consider measurement models of this form in most of our specific examples for simplicity, but two-step estimation is not limited to this case. In the simulation studies Section 4 we also include models where some items measure more than one latent trait.
Specification of the measurement models for individual items $Y_{ijk}$ depends on the forms of the items. We consider situations where each item is a categorical variable with a finite number of possible values. Most of our specific examples will be ones where the items are binary. Coding their values as 0 and 1, in that case we specify the model for each item in (7) as the logit model
@@
so that $\boldsymbol{\theta}_{1k}=(\tau_{k1},\dots,\tau_{kp_{k}},\lambda_{k1},\dots,\lambda_{kp_{k}})^{\prime}$. In the language of item response theory modelling, this is a two-parameter logistic (2-PL) model. There can be parameter constraints within $\boldsymbol{\theta}_{1k}$, such as taking all $\lambda_{kj}$ for the same $k$ to be equal. Other model specifications will be used if an item has $L>2$ categories. If they are regarded as ordered, the binary model can be replaced by, for example, the ordinal logistic model $\text{logit}[P(Y_{ikj}\leq l|\eta_{ik};\boldsymbol{\theta}_{1k})]=\tau_{kjl}-\lambda_{kj}\eta_{ik}$ for categories $l=1,\dots,L-1$. For items with unordered categories, we can use the multinomial logistic model $\log[P(Y_{ikj}=l|\eta_{ik};\boldsymbol{\theta}_{1k})/P(Y_{ikj}=1|\eta_{ik};\boldsymbol{\theta}_{1k})]=\tau_{kjl}+\lambda_{kjl}\eta_{ik}$ for $l=2,\dots,L$.
@@ 3.1 Two-step point estimates
[…]
where the structural and measurement models are specified as described in Section 2. The measurement model on the second line of (9) is for the simple measurement structure where each item measures exactly one latent variable $\eta_{ik}$, and it would be modified appropriately if this was not the case. Maximizing (9) with respect to $\boldsymbol{\theta}=(\boldsymbol{\theta}_{1}^{\prime},\boldsymbol{\theta}_{2}^{\prime})^{\prime}$ would give the overall maximum likelihood estimates (one-step estimates) of $\boldsymbol{\theta}$; we will discuss them further in Section 3.4.1 The integrals in (9) and in (10) below are not available in a closed form, so numerical integration is needed to evaluate them in estimation.
In step 1 of two-step estimation, the measurement parameters $\boldsymbol{\theta}_{1}$ are estimated from a simpler specification. When the measurement structure is simple, this can be done separately for each $\eta_{ik}$, using the log-likelihoods
[…]
@@
Compared with latent class models (as discussed in Bakk and Kuha 2018), there are some additional considerations here where the latent variables $\boldsymbol{\eta}_{i}$ are continuous. First, as usual for any models with continuous latent variables, some parameter constraints are needed to fix the scales of $\boldsymbol{\eta}_{i}$ and thus to identify the parameters $\boldsymbol{\theta}$. In two-step estimation they are imposed in the first step, and the scale implied by them then carries over to the second step where no further constraints are required. We identify the scale of each $\eta_{ik}$ by fixing parameters in the measurement model, for example $\tau_{kj}=0$ and $\lambda_{kj}=1$ for one $j$ in the logistic model in (8). This anchors the latent scale in the items $\mathbf{Y}_{i}$, leaving the distribution of $\boldsymbol{\eta}_{i}$ free. An alternative first-step constraint would be to set $\mu_{k}=0$ and $\sigma^{2}_{k}=1$ for all $k$, implying that each latent variable is on a scale where its marginal distribution is standard normal.
@@
Another difference to typical latent class models is that in applications of latent trait models it is more common for the latent $\boldsymbol{\eta}_{i}$ to be multivariate. The measurement model can then be correspondingly more complex. Our recommendation is that step 1 of the estimation should then be divided into the simplest models that allow correct estimation of the parameters of the specified measurement model. When this model has simple structure, this means estimating single-trait models (10) for each latent trait separately. In other cases this is generalised only as far as is needed. For example, suppose that the only deviation from a simple measurement structure is that some items measure both $\eta_{i1}$ and $\eta_{i2}$. Then the measurement parameters for them would be estimated from a two-trait model with a bivariate normal distribution for ($\eta_{i1},\eta_{i2}$), and the rest of $\tilde{\boldsymbol{\theta}}_{1}$ from one-trait models for any other traits. This keeps step 1 as simple as possible, which is a key benefit of two-step estimation. An alternative would be to estimate all of the measurement models at once, from a $K$-trait model for a $K$-variate normal $\boldsymbol{\eta}_{i}$. In other words, this would be obtained by omitting just $\mathbf{X}_{i}$ from the structural model, and ignoring any parameter constraints in it. This, however, would make the first step more complex, often with no clear benefit. Rosseel and Loh (2022) give a good discussion of the considerations of how to organise step 1 in this respect. They too recommend estimating the measurement model of each distinct block of latent variables separately in most cases.
@@
In practice some of the observed variables $\mathbf{X}_{i}$ and $\mathbf{Y}_{i}$ may be missing. Any missing items in $\mathbf{Y}_{i}$ are taken to be missing at random (MAR), and their contributions are simply omitted from (9) and (10). In (9) we take $\mathbf{X}_{i}$ to be complete, implying that units with any missing values in $\mathbf{X}_{i}$ are omitted from estimation of the structural model. This gives valid estimates for the model as long as the missingness in $\mathbf{X}_{i}$ does not depend on the response variables $\boldsymbol{\eta}_{i}$ (or on $\mathbf{Y}_{i}$). Multiple imputation could be used to include also incomplete observations of $\mathbf{X}_{i}$. These assumptions are conventional, and their implications are mostly the same for both two-step and one-step estimation. There are, however, some small differences which favour the two-step approach here. First, if the MAR assumption fails only for some items, say those in $\mathbf{Y}_{il}$, this affects only the estimates of the measurement model they contribute to, i.e. only $\tilde{\boldsymbol{\theta}}_{1l}$. This is because $\tilde{\boldsymbol{\theta}}_{1k}$ for different $k$ are estimated separately from each other in step 1. Second, even if we omit some units with incomplete data on $\mathbf{X}_{i}$ in step 2, we can still use them in step 1 where $\mathbf{X}_{i}$ are not involved, hence reducing the loss of information from missing data.

### v3

@@ Abstract
We consider likelihood-based two-step estimation of latent variable models, in which just the measurement model is estimated in the first step and the measurement parameters are then fixed at their estimated values in the second step where the structural model is estimated. We show how this approach can be implemented for latent trait models (item response theory models) where the latent variables are continuous and their measurement indicators are categorical variables. The properties of two-step estimators are examined using simulation studies and applied examples. They perform well, and have attractive practical and conceptual properties compared to the alternative one-step and three-step approaches. These results are in line with expectations from theory and with previous findings for other families of latent variable models. This provides strong evidence that two-step estimation is a flexible and useful general method of estimation for different types of latent variable models.
Key words: Item response theory models; generalized latent variable models; structural equation models; pseudo-maximum likelihood estimation
@@ 1 Introduction
General latent variable models have two main parts: a measurement model which describes how the latent variables that appear in the model are measured by observed indicators of them, and a structural model which describes the associations among the latent variables and any observed explanatory and response variables which are not treated as measurement indicators. For instance, in the illustrative example considered in Section 5 of this paper the structural model specifies how individuals’ extrinsic and intrinsic work value orientations are associated with characteristics of the individuals, and the measurement model specifies how these value orientations are measured by a set of survey questions. Here, as in many applications, the structural model is the focus of substantive interest, but the measurement model also needs to be included in order for the structural model to be estimable.
Estimation of these elements can be organised in different ways. In joint or one-step estimation, both parts of the model are estimated together. When this is done by maximizing the joint likelihood, one-step estimates are the maximum likelihood (ML) estimates of the model parameters, also known in this context as full-information ML (FIML) estimates.. In contrast, “stepwise” approaches divide the estimation into separate steps. The most familiar of them is three-step estimation. In its first step, the measurement model is estimated from a specification which omits all or most of the structural model. In the second step, this estimated measurement model is used to assign predicted values of the latent variables to the units of analysis, such as factor scores for continuous latent variables. In the third step, these scores are used in place of the latent variables to estimate the structural model. There are “naive” and “adjusted” versions of three-step estimation, depending on whether it attempts to account for the measurement error that results from using the predicted values.
@@
Two-step estimation is a general idea, and one of the points that we want to emphasise is that it can in principle be applied to any latent variable models. However, there may be differences in performance of the estimates, ease of implementation, or other points to consider when it is used for different broad classes of models. In the literature so far, two-step estimation has accordingly been described separately for different types of models. It was proposed for latent class analysis, that is for models where both the latent variables and their indicators are categorical, by Bandeen-Roche et al. (1997), Xue and Bandeen-Roche (2002) and Bakk and Kuha (2018), and has since been extended to further versions of them (e.g. multilevel latent class analysis in Di Mari et al. 2023). For structural equation models (SEMs), where the latent variables and indicators are both continuous, the idea was introduced already by Burt (1973, 1976), but detailed exploration of it is much more recent. In particular, Rosseel and Loh (2024) proposed a general two-step approach for SEMs, with the title of “structural-after-measurement” (SAM) estimation. They describe two (usually equivalent) forms of it: “local SAM” which summarises the first step in the form of the estimated means and covariance matrix of the variables in the structural model, and “global” SAM which is analogous to the approach that we describe in this paper for latent trait models. A method for SEMs is also proposed by Levy (2023), using Bayesian (MCMC) estimation for both steps, and Levy and McNeish (2025) extended it to the kinds of models with categorical indicators that we will also consider here. Another closely related paper is Skrondal and Kuha (2012), who use two-step estimation to correct for covariate measurement error in regression models. Two-step methods for these classes of models are now also being implemented in general-purpose software for latent variable modelling, in Latent Gold for latent class analysis (Vermunt and Magidson, 2021b) and in the R package lavaan for SAM (Rosseel, 2012; Rosseel and Loh, 2024). A comprehensive recent review of two-step and other stepwise methods is given by Vermunt (2025).
@@
In this paper we propose and examine two-step estimation for another family of latent variable models, one where the latent variables are continuous but their indicators are categorical. These are commonly known as Item Response Theory (IRT) models, and they can also be placed in a more general context as instances of generalised linear latent and mixed models (GLLAMMs; Skrondal and Rabe-Hesketh 2004) or as structural models combined with general linear latent variable measurement models (GLLVMs; Bartholomew et al. 2011). We refer to them as latent trait models. They resemble SEMs in having continuous latent variables, and standard latent class models in having categorical indicators. Like SEMs, they are often used with complex structural models which include multiple latent variables, whereas latent class analysis more often includes only one latent class variable. Computationally, latent trait models are typically the most demanding of these families, at least for ML estimation, because their log likelihood cannot be expressed in a closed form.
@@
The methods of estimation that we consider throughout are likelihood-based. The starting point is thus that one-step FIML estimation would be considered appropriate, but we look for the kinds of further advantages that stepwise modifications of it can offer. The merits of likelihood-based estimation itself are not at issue here. Such a focus is standard when these models are approached from an IRT perspective or as general statistical models with latent variables. Other methods of estimation are also available. In particular, in structural equation modelling literature, where these models would be seen as SEMs with categorical indicators, it is also common to use limited-information methods such as different forms of weighted least squares estimation. This choice is distinct from that of the number of steps in the procedure. In other words, both one-step and two-step methods can also be combined with different methods of estimation, with appropriate modifications to their implementation and theoretical justification. We discuss this topic briefly further in Section 3.6.
The latent trait models that we consider are introduced in Section 2. The definition, implementation and properties of two-step estimation for them are described in Section 3. Simulation studies are reported in Section 4 and the real-data example in Section 5, and concluding discussion is given in Section 6. Extended tables of some of the results from the simulations and data analysis are given in a supplementary appendix. Computer code for the estimation in R and Mplus (both of which are needed) and replication code for the applied example are also provided as supplementary materials.
@@ 2 Models and variables
In two-step estimation, $\boldsymbol{\theta}_{1}$ are estimated first and then fixed at their estimated values in the second step where $\boldsymbol{\theta}_{2}$ are estimated. This approach can be used for any instance of this general model. For specificity, however, in most of this paper we focus on its use for a particular class of latent variable models. Here the items in $\mathbf{Y}_{i}$ are categorical variables, $\boldsymbol{\eta}_{i}$ are continuous and normally distributed, and the measurement model is of a certain common form. This specification is defined in this section. All of its elements can be further relaxed, as discussed in Section 3.7 below.
@@
where $\mathbf{B}_{0}$, $\mathbf{B}_{\eta}$ and $\mathbf{B}_{x}$ are matrices of parameters, and $\boldsymbol{\zeta}_{i}\sim N(\mathbf{0},\boldsymbol{\Psi})$ are normally distributed model residuals. This is a variant of a formulation that is commonly used to specify structural equation models in a compact form. For our purposes, however, it is more convenient to write the model in a different way, in terms of a chain of conditional distributions. Suppose that $\boldsymbol{\eta}_{i}=(\eta_{i1},\dots,\eta_{iK})^{\prime}$ includes $K\geq 1$ variables, and that the model specification further groups them into $1\leq M\leq K$ blocks $\boldsymbol{\eta}_{i(m)}$, so that we can also write $\boldsymbol{\eta}_{i}=(\boldsymbol{\eta}_{i(1)}^{\prime},\dots,\boldsymbol{\eta}_{i(M)}^{\prime})^{\prime}$ in this pre-specified order. We assume that (1) is recursive, i.e. that $\mathbf{B}_{\eta}$ is lower triangular and $\boldsymbol{\Psi}$ is block diagonal in the order of the blocks in $\boldsymbol{\eta}_{i}$. The structural model can then be written as
@@
$\hskip 10.00002pt\times\cdots\times p(\boldsymbol{\eta}_{i(M)}|\mathbf{X}_{i},\boldsymbol{\eta}_{i(1)},\dots,\boldsymbol{\eta}_{i(M-1)};\boldsymbol{\theta}_{2M})$ (2)
where the structural parameters are partitioned correspondingly as $\boldsymbol{\theta}_{2}=(\boldsymbol{\theta}_{21}^{\prime},\dots,\boldsymbol{\theta}_{2M}^{\prime})^{\prime}$. The conditional distributions are all (multivariate or univariate) normal and specified by the linear models
@@
$p(\boldsymbol{\eta}_{i(m)}|\mathbf{X}_{i},\boldsymbol{\eta}_{i(1)},\dots,\boldsymbol{\eta}_{i(m-1)};\boldsymbol{\theta}_{2m})$ $\sim$ $\hskip-5.0ptN(\boldsymbol{\beta}_{m0}+\boldsymbol{\beta}_{mx}\mathbf{X}_{i}+\sum_{l=1}^{m-1}\boldsymbol{\beta}_{ml}\boldsymbol{\eta}_{i(l)},\,\boldsymbol{\Psi}_{m})$ (4)
for $m=2,\dots,M$. Each of $\boldsymbol{\theta}_{21},\dots,\boldsymbol{\theta}_{2M}$ thus consists of the corresponding $\boldsymbol{\beta}$ and $\boldsymbol{\Psi}$ parameters. Constraints on elements of $\boldsymbol{\theta}_{2}$ can be included, most often zero constraints on the $\boldsymbol{\beta}$ parameters which omit some of the explanatory variables from some parts of the model and thus yield different non-saturated choices for the overall structural model (2). If observed covariates $\mathbf{X}_{i}$ are not present, they are omitted from (2)–(4). If any $\boldsymbol{\eta}_{i(m)}$ are actually observed (rather than latent) variables, it would also be straightfoward to replace the linear model with some other model for them. For example, if $\boldsymbol{\eta}_{i(M)}$ is a single binary observed variable, we could specify a binary logistic model for it given $\mathbf{X}_{i}$ and $\boldsymbol{\eta}_{i(1)},\dots,\boldsymbol{\eta}_{i(M-1)}$.
Consider then the measurement model. We assume that $p(\mathbf{Y}_{i}|\boldsymbol{\eta}_{i},\mathbf{X}_{i};\boldsymbol{\theta}_{1})=p(\mathbf{Y}_{i}|\boldsymbol{\eta}_{i};\boldsymbol{\theta}_{1})$, so that there is no non-equivalence of measurement (differential item functioning) with respect to observed covariates $\mathbf{X}_{i}$. The overall model can then be written as
@@
$p(\mathbf{Y}_{i}|\mathbf{X}_{i};\boldsymbol{\theta})=\int\,p(\mathbf{Y}_{i}|\boldsymbol{\eta}_{i};\boldsymbol{\theta}_{1})\,p(\boldsymbol{\eta}_{i}|\mathbf{X}_{i};\boldsymbol{\theta}_{2})\,d\boldsymbol{\eta}_{i}.$ (6)
@@
so that $\boldsymbol{\theta}_{1}=(\boldsymbol{\theta}_{11}^{\prime},\dots,\boldsymbol{\theta}_{1K}^{\prime})^{\prime}$, with the different $\boldsymbol{\theta}_{1k}$ taken to be distinct from each other. We consider measurement models of this form in most of our specific examples for simplicity, but two-step estimation is not limited to this case (in the simulation studies Section 4 we also include models where some items measure more than one latent trait).
Specification of the measurement models for individual items $Y_{ijk}$ depends on the forms of the items. We consider situations where each item is a categorical variable with a finite number of possible values. In most of our specific examples the items are binary. Coding their values as 0 and 1, in that case we specify the model for each item in (7) as the logit model
@@
so that $\boldsymbol{\theta}_{1k}=(\tau_{k1},\dots,\tau_{kp_{k}},\lambda_{k1},\dots,\lambda_{kp_{k}})^{\prime}$. In the language of IRT modelling, this is a two-parameter logistic (2-PL) model. There can be parameter constraints within $\boldsymbol{\theta}_{1k}$, such as taking all $\lambda_{kj}$ for the same $k$ to be equal. Other model specifications will be used if an item has $L>2$ categories. If they are regarded as ordered, the binary model can be replaced by, for example, the ordinal logistic model $\text{logit}[P(Y_{ikj}\leq l|\eta_{ik};\boldsymbol{\theta}_{1k})]=\tau_{kjl}-\lambda_{kj}\eta_{ik}$ for categories $l=1,\dots,L-1$. For items with unordered categories, we can use the multinomial logistic model $\log[P(Y_{ikj}=l|\eta_{ik};\boldsymbol{\theta}_{1k})/P(Y_{ikj}=1|\eta_{ik};\boldsymbol{\theta}_{1k})]=\tau_{kjl}+\lambda_{kjl}\eta_{ik}$ for $l=2,\dots,L$.
@@ 3.1 Two-step point estimates
$\ell(\boldsymbol{\theta})=\sum_{i=1}^{n}\,\log p(\mathbf{Y}_{i}|\mathbf{X}_{i};\boldsymbol{\theta})=\sum_{i=1}^{n}\,\log\,\int p(\mathbf{Y}_{i}|\boldsymbol{\eta}_{i};\boldsymbol{\theta}_{1})\,p(\boldsymbol{\eta}_{i}|\mathbf{X}_{i};\boldsymbol{\theta}_{2})\,d\boldsymbol{\eta}_{i}$ (9)
$=$ $\sum_{i=1}^{n}\,\log\int\,\left[\prod_{k=1}^{K}\prod_{j=1}^{p_{k}}p(Y_{ikj}|\eta_{ik};\boldsymbol{\theta}_{1k})\right]\,p(\boldsymbol{\eta}_{i}|\mathbf{X}_{i};\boldsymbol{\theta}_{2})\,d\boldsymbol{\eta}_{i}$
where the structural and measurement models are specified as described in Section 2. The measurement model on the second line of (9) is for the simple measurement structure where each item measures exactly one latent variable $\eta_{ik}$, and it would be modified appropriately if this was not the case. Maximizing (9) with respect to $\boldsymbol{\theta}=(\boldsymbol{\theta}_{1}^{\prime},\boldsymbol{\theta}_{2}^{\prime})^{\prime}$ would give the overall (full-information) maximum likelihood estimates of $\boldsymbol{\theta}$, i.e. the one-step estimates.
In step 1 of two-step estimation, the measurement parameters $\boldsymbol{\theta}_{1}$ are estimated. This can be done using any specification that allows consistent estimation of $\boldsymbol{\theta}_{1}$. The default would be to use the simplest model with that property. When the measurement structure is simple, this means that the step-1 estimation can be done separately for each $\eta_{ik}$, using the log-likelihoods
$\ell(\boldsymbol{\psi}_{k})=\sum_{i=1}^{n}\,\log\int\left[\prod_{j=1}^{p_{k}}\,p(Y_{ikj}|\eta_{ik};\boldsymbol{\theta}_{1k})\right]\,p(\eta_{ik};\mu_{k},\sigma^{2}_{k})\,d\eta_{ik}$ (10)
@@
Compared with latent class models (as discussed in Bakk and Kuha 2018), there are some additional considerations here where the latent variables $\boldsymbol{\eta}_{i}$ are continuous. First, some parameter constraints are needed to fix the scales of $\boldsymbol{\eta}_{i}$ and identify the parameters $\boldsymbol{\theta}$. In two-step estimation they are imposed in the step 1, and the scale implied by them then carries over to the step 2 where no further constraints are required. We identify the scale of each $\eta_{ik}$ by fixing parameters in the measurement model, for example $\tau_{kj}=0$ and $\lambda_{kj}=1$ for one $j$ in the logistic model in (8). This anchors the latent scale in the items $\mathbf{Y}_{i}$, leaving the distribution of $\boldsymbol{\eta}_{i}$ free. An alternative first-step constraint would be to set $\mu_{k}=0$ and $\sigma^{2}_{k}=1$ for all $k$, implying that each latent variable is on a scale where its marginal distribution is standard normal.
@@
Another difference to latent class models is that in applications of latent trait models it is more common for the latent $\boldsymbol{\eta}_{i}$ to be multivariate. The measurement model can then be correspondingly more complex. When it has simple structure, we recomment that step-1 estimation is carried out for the single-trait models (10) for each latent trait separately. In other cases this is generalised only as far as is needed. For example, suppose that the only deviation from a simple measurement structure is that some items measure both $\eta_{i1}$ and $\eta_{i2}$. Then the measurement parameters for them would be estimated from a two-trait model with a bivariate normal distribution for ($\eta_{i1},\eta_{i2}$), and the rest of $\boldsymbol{\theta}_{1}$ from one-trait models for any other traits. An alternative would be to estimate the whole measurement model at once, from a $K$-trait model for a $K$-variate normal $\boldsymbol{\eta}_{i}$. In other words, this would be obtained by omitting just $\mathbf{X}_{i}$ from the structural model, and ignoring any parameter constraints in it. This, however, would make the first step more complex, with no clear benefit. Rosseel and Loh (2024) give a good discussion of the considerations of how to organise step 1 in this respect. They too recommend estimating the measurement model of each distinct block of latent variables separately in most cases.
@@
In practice some of the observed variables $\mathbf{X}_{i}$ and $\mathbf{Y}_{i}$ may be missing. Any missing items in $\mathbf{Y}_{i}$ are taken to be missing at random (MAR), and their contributions are simply omitted from (9) and (10). In (9) we take $\mathbf{X}_{i}$ to be complete, implying that units with any missing values in $\mathbf{X}_{i}$ are omitted from estimation of the structural model. This gives valid estimates for the model as long as the missingness in $\mathbf{X}_{i}$ does not depend on the response variables $\boldsymbol{\eta}_{i}$ (or on $\mathbf{Y}_{i}$). Multiple imputation could be used to include also incomplete observations of $\mathbf{X}_{i}$. These assumptions are conventional, and their implications are mostly the same for both two-step and one-step estimation. There are, however, some small differences which favour the two-step approach here. First, if the MAR assumption fails only for some items, say those in $\mathbf{Y}_{il}$, this affects only the estimates of the measurement model they contribute to, i.e. only $\tilde{\boldsymbol{\theta}}_{1l}$. Second, even if we omit some units with incomplete data on $\mathbf{X}_{i}$ in step 2, we can still use them in step 1 where $\mathbf{X}_{i}$ are not involved, hence reducing the loss of information from missing data.